OpenStudy (anonymous):

Need Help! How do you solve for (3/7)^7/3 to be in fraction form?

OpenStudy (mathmate):

Is the question \(\Large \frac{(\frac{3}{7})^7}{3}\) or is it \(\Large (\frac{3}{7})^\frac{7}{3}\) ? In any case, use \((\frac{A}{B})^X=\frac{A^X}{B^X}\) and \(M^\frac{x}{y}=\sqrt[y]{M^x}\) to help you.

OpenStudy (anonymous):

It's the second one!

OpenStudy (anonymous):

So if I were using the first method, it would be something like: 3^7/3 all over 7^7/3 ?

OpenStudy (anonymous):

But what would I do after that?

OpenStudy (mathmate):

@mary_am Please remember the priority of operations require that exponentiation is before multiplication and division, so use brackets as required. \(\Large (\frac{3}{7})^\frac{7}{3}\) is written (3/7)^(7/3) to avoid further confusion.

OpenStudy (mathmate):

"So if I were using the first method, it would be something like: 3^7/3 all over 7^7/3 ? " probably meant = ( 3^(7/3) ) / ( 7^(7/3) ) Yes written that way would be correct.

OpenStudy (anonymous):

Okay, and then what would happen?

OpenStudy (mathmate):

Then you would use the second rule to make it a simple fraction.

OpenStudy (anonymous):